OpenStudy (anonymous):

derive y= -x^2 (1+x^2-x^3 ) using the definition of derivative? please help me..

OpenStudy (anonymous):

@mukushla

OpenStudy (anonymous):

Two ways: one is to multiply it out and then use the power rule, the other is to use the product rule. Do you know the product rule?

OpenStudy (anonymous):

Let u = -x^2 v = (1+x^2-x^3 ) what are du and dv?

OpenStudy (anonymous):

Just for reference, the product rule is this: d/dx (uv) = u dv/dx + v du/dx or more simply (uv)' = u v' + v u'

OpenStudy (cwrw238):

can you find u' if u = -x^2 ?

OpenStudy (anonymous):

\(f(x)=-x^2-x^4+x^5 \) by definition \[f'(x)=\lim_{\Delta x \rightarrow 0} \frac{f(x+\Delta x)-f(x)}{\Delta x}\\=\lim_{\Delta x \rightarrow 0} \frac{-(x+\Delta x)^2-(x+\Delta x)^4+(x+\Delta x)^5+x^2+x^4-x^5}{\Delta x}\]

OpenStudy (anonymous):

simplifying and neglecting the terms which consists exponents of \(\Delta x\) with powers greater than or equal to 2 u will have... \[f'(x)=\lim_{\Delta x \rightarrow 0} \frac{-2x\Delta x-4x^3 \Delta x+5x^4\Delta x}{\Delta x}=-2x-4x^3+5x^4\]